A water contains 1 mg/L of PO43, how many mg/L of AlB+ would be needed to lower the P043 concentration to 10-s mg/L? What is the final concentration of Al at equilibrium? Known: AlPO4 (s)AlPO3 pKs 20 *Assume this is the only relevant reaction
Answer #1

10-5 mg PO43- = (10-5 X 10-3/94.97) moles = 1.053X 10-10 moles  PO43-

1mg PO43- = (10-3 /94.97) moles = 1.053X 10-5 moles PO43-

1mole Al3+ reacts with 1 mole PO43-

Let x moles Al3+ reacts with x mole  PO43-

(1.053X 10-5 -x) = 1.053X 10-10

x= 1.053X 10-5 moles Al3+ = 28.4099x 10-5 g Al3+ = 0.284 mg of Al3+

So, the Al3+ required is 0.284 mg/L.

Let the equlibrium concentration is 'S' moles of Al3+,[Al3+] = [ PO43-] = S

Solubility product Ksp= 10-20 = [Al3+] [ PO43-] = S2

S= 10-10 moles of Al3+ remain at equilibrium.